OpenStudy (alibaby):

Find an irrational number that is between 9.5 and 9.7. Explain why it is irrational. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (alibaby):

@zepdrix

zepdrix (zepdrix):

Hmm

zepdrix (zepdrix):

Ok this might be a little confusing... We'll try it though.

zepdrix (zepdrix):

If we square our 9.5 and our 9.7, \(\large\rm 9.5^2=90.25\) \(\large\rm 9.7^2=94.09\) We can choose the square root of a number between these two values and it will be between 9.5 and 9.7. How bout we go with 91. So we know that \(\large\rm \sqrt{91}\) must be between 9.5 and 9.7. When we punch that value into our calculator we get 9.539392014...

zepdrix (zepdrix):

So our irrational number that we came up with is \(\large\rm \sqrt{91}\), that's the exact number. The decimal approximation, to the nearest hundredth will be \(\large\rm 9.54\). I rounded the 3 up to a 4 because of the 9 next to it.

zepdrix (zepdrix):

Confusing with all the squares and stuff? What do you think? :d

OpenStudy (alibaby):

it's not alicia btw it here sis and no not confusing at all

OpenStudy (alibaby):

thanks

zepdrix (zepdrix):

:p

OpenStudy (alibaby):

alicia say skype her

OpenStudy (alibaby):

@zepdrix